# Rigorous dense graph limit of a model for

### v1

@@ 6. The energy functionals
$\int_{\mathcal{I}}\sigma(x)\frac{\delta^{2}p[b;\Phi]}{\delta b^{2}}(x)\,\mathrm{d}x$ $=$ $\frac{1}{2}\int_{\mathcal{I}}\int_{\mathcal{I}}b(x,y)\left(\mathfrak{d}{\frac{\delta p}{\delta b}}(x,y)\right)^{2}\,\mathrm{d}w(x,y)$
$\quad+\frac{1}{2}\int_{\mathcal{I}}\int_{\mathcal{I}}b(x,y)\left(\mathfrak{d}{\frac{\delta^{2}p}{\delta b^{2}}}(x,y)\right)\mathfrak{d}{p}(x,y)\,\mathrm{d}w(x,y)$
$\qquad+\frac{1}{2}\int_{\mathcal{I}}\int_{\mathcal{I}}{\Phi(x,y)}\left(\mathfrak{d}{\frac{\delta p}{\delta b}}(x,y)\right)\mathfrak{d}{p}(x,y)\,\mathrm{d}w(x,y),$
where here and in the sequel we use the short-hand notation $\frac{\delta p}{\delta b}$ for $\frac{\delta p[b;\Phi]}{\delta b}$. The first-order variation of the weak formulation of (4.5) in direction $\Phi\in L^{2}_{+}(\mathcal{I}\times\mathcal{I})$ with test function $\varphi\in L^{2}(\mathcal{I})$ gives
$\int_{\mathcal{I}}\int_{\mathcal{I}}b(x,y)\left(\mathfrak{d}{\frac{\delta p}{\delta b}}(x,y)\right)\mathfrak{d}{\varphi}(x,y)+\Phi(x,y)\mathfrak{d}{p}(x,y)\mathfrak{d}\varphi(x,y)\,\mathrm{d}w(x,y)=0.$ (6.4)
Setting $\varphi:=\frac{\delta p}{\delta b}$, we obtain the identity
$\int_{\mathcal{I}}\int_{\mathcal{I}}\Phi(x,y)\left(\mathfrak{d}{\frac{\delta p}{\delta b}}(x,y)\right)\mathfrak{d}p(x,y)\,\mathrm{d}w(x,y)=-\int_{\mathcal{I}}\int_{\mathcal{I}}b(x,y)\left(\mathfrak{d}{\frac{\delta p}{\delta b}}(x,y)\right)^{2}\,\mathrm{d}w(x,y).$
We now again take a variation of (6.4) in direction $\Phi\in L^{2}_{+}(\mathcal{I}\times\mathcal{I})$ and use the pressure $p=p[b]$ as a test function, which leads to
$2\int_{\mathcal{I}}\int_{\mathcal{I}}\Phi(x,y)\left(\mathfrak{d}{\frac{\delta p}{\delta b}}(x,y)\right)\mathfrak{d}p(x,y)\,\mathrm{d}w(x,y)=-\int_{\mathcal{I}}\int_{\mathcal{I}}b(x,y)\left(\mathfrak{d}{\frac{\delta^{2}p}{\delta b^{2}}}(x,y)\right)\mathfrak{d}p(x,y)\,\mathrm{d}w(x,y).$
Combining the above identities, we arrive at
$\int_{\mathcal{I}}\sigma(x)\frac{\delta^{2}p[b;\Phi]}{\delta c^{2}}(x)\,\mathrm{d}x=\int_{\mathcal{I}}\int_{\mathcal{I}}b(x,y)\left(\mathfrak{d}{\frac{\delta p}{\delta b}}(x,y)\right)^{2}\,\mathrm{d}w(x,y)\geq 0,$
which establishes the sought-for convexity.
@@
[…]
@@
$\mathcal{F}_{\mathrm{kin}}[\mathcal{Z}^{N}[b^{N}]]=\frac{1}{2}\int_{\mathcal{I}}\int_{\mathcal{I}}\mathcal{Z}^{N}[b^{N}]\left(\mathfrak{d}p^{N}\right)^{2}\,\mathrm{d}w(x,y)=\int_{\mathcal{I}}\sigma p^{N}\,\mathrm{d}x,$
@@
$\mathcal{F}_{\mathrm{kin}}^{N}[\mathcal{Z}^{N}[b^{N}]]=\frac{1}{2}\int_{\mathcal{I}}\int_{\mathcal{I}}\mathcal{Z}^{N}[b^{N}]\left(\mathfrak{d}p^{N}_{h}\right)^{2}\,\mathrm{d}w^{N}(x,y)=\int_{\mathcal{I}}\sigma p_{h}^{N}\,\mathrm{d}x,$
@@
$\mathcal{F}_{\mathrm{met}}^{N}[b^{N}]$ $=$ $\int_{\mathcal{I}}\int_{\mathcal{I}}\left(b^{N}(x,y)\right)^{\gamma}\left(\ell^{N}(x,y)\right)^{\gamma+1}\,\mathrm{d}w^{N}(x,y)=\left\|a^{N}\right\|^{\gamma}_{L^{\gamma}(\mathcal{I}\times\mathcal{I})},$
@@
$\left\|a\right\|^{\gamma}_{L^{\gamma}(\mathcal{I}\times\mathcal{I})}\leq\liminf_{N\to\infty}\left\|a^{N}\right\|^{\gamma}_{L^{\gamma}(\mathcal{I}\times\mathcal{I})}=\liminf_{N\to\infty}\mathcal{F}_{\mathrm{met}}^{N}[b^{N}].$ (7.4)
@@
$\mathcal{F}_{\mathrm{met}}[b]=\left\|a\right\|^{\gamma}_{L^{\gamma}(\mathcal{I}\times\mathcal{I})}\leq\liminf_{N\to\infty}\mathcal{F}_{\mathrm{met}}^{N}[b^{N}].$
Combining the claims of Lemmas 7.1 and 7.2, we conclude that for any sequence $b^{N}\in L^{\omega}(\mathcal{I}\times\mathcal{I})$, $\omega:=\max\{2,\gamma\}$, converging weakly in $L^{\omega}(\mathcal{I}\times\mathcal{I})$ to $b\in L^{\omega}(\mathcal{I}\times\mathcal{I})$, we have
@@
$\mathcal{F}_{\mathrm{kin}}[b]$ $=$ $\lim_{N\to\infty}\left[\mathcal{F}_{\mathrm{kin}}[b^{N}]+\int_{\mathcal{I}}(\sigma-\mathcal{Z}^{N}[\sigma])p^{N}\,\mathrm{d}x\right]$
$=$ $\lim_{N\to\infty}\mathcal{F}_{\mathrm{kin}}[b^{N}].$
@@
$F^{N}[B^{N}]\leq F^{N}[r]=\frac{1}{2N^{2}}\sum_{i=1}^{N}\sum_{j=1}^{N}\left((P_{j}-P_{i})^{2}r+\frac{\nu}{\gamma}r^{\gamma}(L^{N}_{ij})^{\gamma+1}\right)W_{ij}^{N},$ (8.1)
where $(P_{i})_{i=1}^{N}$ is a solution of the Kirchhoff law (4.2) with conductivities $r>0$. Since $L^{N}_{ij}\leq 1$, $W^{N}_{ij}\in\{0,1\}$, we have
$\frac{1}{2N^{2}}\sum_{i=1}^{N}\sum_{j=1}^{N}\frac{\nu}{\gamma}r^{\gamma}(L^{N}_{ij})^{\gamma+1}W_{ij}^{N}\leq\frac{\nu}{2\gamma}r^{\gamma}.$
@@
Consequently,, the right-hand side in (8.1) is uniformly bounded. Denoting $b^{N}:=\mathcal{Q}^{N}[B^{N}]$, Lemma 6.1 gives $\mathcal{F}^{N}[b^{N}]=F^{N}[B^{N}]$, and we have
@@
Moreover, we obviously have
$(B^{N}_{ij})^{\gamma}(L^{N}_{ij})^{\gamma+1}W_{ij}^{N}\leq\frac{\gamma}{\nu}F^{N}_{\mathrm{met}}[B^{N}]\qquad\mbox{for all }i,j\in[N],\;N\in\mathbb{N}.$
Since $w^{N}=\mathcal{Q}^{N}[W^{N}]$ and $\ell^{N}=\mathcal{Q}^{N}[L^{N}]$, this implies a uniform bound on the sequence $a^{N}:=b^{N}\left(\ell^{N}\right)^{\frac{\gamma+1}{\gamma}}w^{N}$ in $L^{\infty}(\mathcal{I}\times\mathcal{I})$. Similarly as in the proof of Lemma 7.2, we have
@@
$\left\|a\right\|^{\gamma}_{L^{\gamma}(\mathcal{I}\times\mathcal{I})}\leq\liminf_{N\to\infty}\left\|a^{N}\right\|^{\gamma}_{L^{\gamma}(\mathcal{I}\times\mathcal{I})}=\liminf_{N\to\infty}\mathcal{F}_{\mathrm{met}}^{N}[b^{N}],$
where $a\in L^{\infty}(\mathcal{I}\times\mathcal{I})$ is the weak-* limit (of a subsequence) of $a^{N}$. We identify the limit in the same way as in the proof of Lemma 7.2, using (7). However, now we use the uniform boundedness of $b^{N}w^{N}$ in $L^{2}(\mathcal{I}\times\mathcal{I})$, which follows from the uniform boundedness of $\left(\ell^{N}\right)^{-\frac{\gamma+1}{\gamma}}$ in $L^{2}(\mathcal{I}\times\mathcal{I})$, imposed by assumption (3.6). We conclude that $a=b\ell^{\frac{\gamma+1}{\gamma}}w$, where $b\in L^{2}(\mathcal{I}\times\mathcal{I})$ is the weak limit of $b^{N}$. Therefore,
@@
$\mathcal{F}_{\mathrm{met}}[b]=\left\|a\right\|^{\gamma}_{L^{\gamma}(\mathcal{I}\times\mathcal{I})}\leq\liminf_{N\to\infty}\mathcal{F}_{\mathrm{met}}^{N}[b^{N}].$ (8.2)
@@
and combined with (8.2) we thus have
@@
We claim that $b$ is a global minimizer of the continuum energy functional (4.4)–(4.5) in $L^{\omega}_{r}(\mathcal{I}\times\mathcal{I})$, $\omega=\max\{2,\gamma\}$. For contradiction, let us assume that there exists $\tilde{b}\in L^{\omega}_{r}(\mathcal{I}\times\mathcal{I})$ such that
@@
Then Theorem 4.1 asserts that the sequence $\tilde{b}^{N}:=\mathcal{Z}^{N}[\tilde{b}]$ converges to $\tilde{b}$ in the norm topology of $L^{\omega}(\mathcal{I}\times\mathcal{I})$, and (4.10) gives
@@
where $\tilde{B}^{N}\in\mathbb{B}^{N}_{r}$ is the matrix of the values of $\tilde{b}^{N}$, i.e., $\tilde{b}^{N}=\mathcal{Q}^{N}[B^{N}]$. Therefore,
$\mathcal{F}[\tilde{b}]=\lim_{N\to\infty}\mathcal{F}^{N}[\tilde{b}^{N}]\geq\liminf_{N\to\infty}\mathcal{F}^{N}[b^{N}]\geq\mathcal{F}[b],$

### v2

@@ 6. The energy functionals
$\mathcal{F}_{\mathrm{kin}}[b]=-2\min_{v\in L^{2}_{0}(\mathcal{I})}\mathcal{J}[v;b]=\sup_{v\in L^{2}_{0}(\mathcal{I})}\left\{2\int_{\mathcal{I}}\sigma v\,\,\mathrm{d}x-\frac{1}{2}\iint_{\mathcal{I}\times\mathcal{I}}b(\overline{\nabla}v)^{2}\,\,\mathrm{d}w\right\}.$ (6.4)
For each fixed $v\in L^{2}_{0}(\mathcal{I})$ the map
$b\ \mapsto\ 2\int_{\mathcal{I}}\sigma(x)v(x)\,\,\mathrm{d}x-\frac{1}{2}\int_{\mathcal{I}}\int_{\mathcal{I}}b(x,y)(\overline{\nabla}v(x,y))^{2}\,\,\mathrm{d}w(x,y)$
is affine in $b$. Therefore (6.4) expresses $\mathcal{F}_{\mathrm{kin}}$ as a supremum of affine functions of $b$, which implies that $\mathcal{F}_{\mathrm{kin}}$ is convex on $L_{r}^{\infty}(\mathcal{I}\times\mathcal{I})$.
Now, let $b_{0},b_{1}\in L^{2}_{r}(\mathcal{I}\times\mathcal{I})$ and $\theta\in[0,1]$. For $M\geq r$ define the truncations
$b_{i}^{M}:=\min\{b_{i},M\}\in L_{r}^{\infty}(\mathcal{I}\times\mathcal{I}),\qquad i\in\{0,1\}.$
Set $b_{\theta}:=(1-\theta)b_{0}+\theta b_{1}$ and $b_{\theta}^{M}:=(1-\theta)b_{0}^{M}+\theta b_{1}^{M}$. We have, for every $M\geq r$,
$\mathcal{F}_{\mathrm{kin}}[b_{\theta}^{M}]\ \leq\ (1-\theta)\mathcal{F}_{\mathrm{kin}}[b_{0}^{M}]+\theta\mathcal{F}_{\mathrm{kin}}[b_{1}^{M}].$
Moreover, $b_{i}^{M}\to b_{i}$ strongly in $L^{2}(\mathcal{I}\times\mathcal{I})$ as $M\to\infty$, hence also $b_{\theta}^{M}\to b_{\theta}$ strongly in $L^{2}(\mathcal{I}\times\mathcal{I})$. By the strong continuity of $\mathcal{F}_{\mathrm{kin}}$ on $L^{2}_{r}(\mathcal{I}\times\mathcal{I})$ established in Lemma 6.3 we conclude
$\mathcal{F}_{\mathrm{kin}}[b_{\theta}]$ $=\lim_{M\to\infty}\mathcal{F}_{\mathrm{kin}}[b_{\theta}^{M}]$
$\leq(1-\theta)\lim_{M\to\infty}\mathcal{F}_{\mathrm{kin}}[b_{0}^{M}]+\theta\lim_{M\to\infty}\mathcal{F}_{\mathrm{kin}}[b_{1}^{M}]$
$=(1-\theta)\mathcal{F}_{\mathrm{kin}}[b_{0}]+\theta\mathcal{F}_{\mathrm{kin}}[b_{1}].$
This proves convexity of $\mathcal{F}_{\mathrm{kin}}$ on $L^{2}_{r}(\mathcal{I}\times\mathcal{I})$.
@@
We organize the proof into the following four lemmas, working separately with the kinetic and metabolic parts of the energy functional. We start by proving the $\liminf$-part of the $\Gamma$-convergence, i.e., claim (4.9) of Theorem 4.1.
@@
$\mathcal{F}_{\mathrm{kin}}[\mathcal{Z}^{N}[b^{N}]]=\frac{1}{2}\int_{\mathcal{I}}\int_{\mathcal{I}}\mathcal{Z}^{N}[b^{N}]\left(\overline{\nabla}p^{N}\right)^{2}\,\mathrm{d}w(x,y)=\int_{\mathcal{I}}\sigma p^{N}\,\mathrm{d}x,$
@@
$\mathcal{F}_{\mathrm{kin}}^{N}[\mathcal{Z}^{N}[b^{N}]]=\frac{1}{2}\int_{\mathcal{I}}\int_{\mathcal{I}}\mathcal{Z}^{N}[b^{N}]\left(\overline{\nabla}p^{N}_{h}\right)^{2}\,\mathrm{d}w^{N}(x,y)=\int_{\mathcal{I}}\sigma p_{h}^{N}\,\mathrm{d}x,$
@@
$\mathcal{F}_{\mathrm{met}}^{N}[b^{N}]$ $=$ $\frac{\nu}{2\gamma}\int_{\mathcal{I}}\int_{\mathcal{I}}\left(b^{N}(x,y)\right)^{\gamma}\left(\ell^{N}(x,y)\right)^{\gamma+1}\,\mathrm{d}w^{N}(x,y)=\frac{\nu}{2\gamma}\left\|a^{N}\right\|^{\gamma}_{L^{\gamma}(\mathcal{I}\times\mathcal{I})},$
@@
$\left\|a\right\|^{\gamma}_{L^{\gamma}(\mathcal{I}\times\mathcal{I})}\leq\liminf_{N\to\infty}\left\|a^{N}\right\|^{\gamma}_{L^{\gamma}(\mathcal{I}\times\mathcal{I})}=\frac{2\gamma}{\nu}\liminf_{N\to\infty}\mathcal{F}_{\mathrm{met}}^{N}[b^{N}].$ (7.4)
@@
$\mathcal{F}_{\mathrm{met}}[b]=\frac{\nu}{2\gamma}\left\|a\right\|^{\gamma}_{L^{\gamma}(\mathcal{I}\times\mathcal{I})}\leq\liminf_{N\to\infty}\mathcal{F}_{\mathrm{met}}^{N}[b^{N}].$
Combining the claims of Lemmas 7.1 and 7.2, we conclude that for any sequence $b^{N}\in L_{r}^{\omega}(\mathcal{I}\times\mathcal{I})$, $\omega:=\max\{2,\gamma\}$, converging weakly in $L^{\omega}(\mathcal{I}\times\mathcal{I})$ to $b\in L^{\omega}(\mathcal{I}\times\mathcal{I})$, we have
@@
$\mathcal{F}_{\mathrm{kin}}[b]$ $=$ $\lim_{N\to\infty}\left[\mathcal{F}_{\mathrm{kin}}^{N}[b^{N}]+\int_{\mathcal{I}}(\sigma-\mathcal{Z}^{N}[\sigma])p^{N}\,\mathrm{d}x\right]$
$=$ $\lim_{N\to\infty}\mathcal{F}_{\mathrm{kin}}^{N}[b^{N}].$
@@
$F^{N}[B^{N}]\leq F^{N}[r]=\frac{1}{2N^{2}}\sum_{i=1}^{N}\sum_{j=1}^{N}\left((P_{j}-P_{i})^{2}r+\frac{\nu}{\gamma}r^{\gamma}(\mathrm{L}^{N}_{ij})^{\gamma+1}\right)W_{ij}^{N},$ (8.1)
where $(P_{i})_{i=1}^{N}$ is a solution of the Kirchhoff law (4.2) with conductivities $r>0$. Since $\mathrm{L}^{N}_{ij}\leq 1$ and $W^{N}_{ij}\in\{0,1\}$, we have
$\frac{1}{2N^{2}}\sum_{i=1}^{N}\sum_{j=1}^{N}\frac{\nu}{\gamma}r^{\gamma}(\mathrm{L}^{N}_{ij})^{\gamma+1}W_{ij}^{N}\leq\frac{\nu}{2\gamma}r^{\gamma}.$
@@
Consequently, the right-hand side in (8.1) is uniformly bounded. Denoting $b^{N}:=\mathcal{Q}^{N}[B^{N}]$, Lemma 6.1 gives $\mathcal{F}^{N}[b^{N}]=F^{N}[B^{N}]$, and we have
@@
We define the sequence $a^{N}:=b^{N}\left(\ell^{N}\right)^{\frac{\gamma+1}{\gamma}}w^{N}$. Note that the metabolic energy can be written as
$\mathcal{F}_{\mathrm{met}}^{N}[b^{N}]=\frac{\nu}{2\gamma}\int_{\mathcal{I}}\int_{\mathcal{I}}\left(b^{N}\right)^{\gamma}\left(\ell^{N}\right)^{\gamma+1}\,\mathrm{d}w^{N}(x,y)=\frac{\nu}{2\gamma}\left\|a^{N}\right\|^{\gamma}_{L^{\gamma}(\mathcal{I}\times\mathcal{I})}.$
The uniform bound on $\mathcal{F}_{\mathrm{met}}^{N}[b^{N}]$ implies that the sequence $a^{N}$ is uniformly bounded in $L^{\gamma}(\mathcal{I}\times\mathcal{I})$. Since $\gamma>1$, the space $L^{\gamma}(\mathcal{I}\times\mathcal{I})$ is reflexive, and there exists a subsequence (not relabeled) such that $a^{N}\rightharpoonup a$ weakly in $L^{\gamma}(\mathcal{I}\times\mathcal{I})$. Using the weak lower semicontinuity of the $L^{\gamma}$-norm, we have
@@
$\left\|a\right\|^{\gamma}_{L^{\gamma}(\mathcal{I}\times\mathcal{I})}\leq\liminf_{N\to\infty}\left\|a^{N}\right\|^{\gamma}_{L^{\gamma}(\mathcal{I}\times\mathcal{I})}=\frac{2\gamma}{\nu}\liminf_{N\to\infty}\mathcal{F}_{\mathrm{met}}^{N}[b^{N}].$
We identify the limit $a\in L^{\gamma}(\mathcal{I}\times\mathcal{I})$. Since $b^{N}=\mathcal{Q}^{N}[B^{N}]$ and $B^{N}_{ij}=0$ whenever $W^{N}_{ij}=0$, we have $b^{N}=b^{N}w^{N}$, and thus $a^{N}=b^{N}\left(\ell^{N}\right)^{\frac{\gamma+1}{\gamma}}$. Consequently,
$b^{N}=a^{N}\left(\ell^{N}\right)^{-\frac{\gamma+1}{\gamma}}.$
Since $\gamma>2$, by Hölder’s inequality we have
$\left\|b^{N}\right\|_{L^{2}(\mathcal{I}\times\mathcal{I})}\leq\left\|a^{N}\right\|_{L^{\gamma}(\mathcal{I}\times\mathcal{I})}\left\|\left(\ell^{N}\right)^{-\frac{\gamma+1}{\gamma}}\right\|_{L^{\frac{2\gamma}{\gamma-2}}(\mathcal{I}\times\mathcal{I})}=\left\|a^{N}\right\|_{L^{\gamma}(\mathcal{I}\times\mathcal{I})}\left\|\left(\ell^{N}\right)^{-1}\right\|_{L^{\frac{2(\gamma+1)}{\gamma-2}}(\mathcal{I}\times\mathcal{I})}^{\frac{\gamma+1}{\gamma}},$
and the right-hand side is uniformly bounded by assumption (3.6). Therefore, there exists $b\in L^{2}(\mathcal{I}\times\mathcal{I})$ and a subsequence (not relabeled) such that $b^{N}\rightharpoonup b$ weakly in $L^{2}(\mathcal{I}\times\mathcal{I})$.
Now, for any $\varphi\in L^{\infty}(\mathcal{I}\times\mathcal{I})$ we have
$\int_{\mathcal{I}}\int_{\mathcal{I}}a^{N}\varphi\,\mathrm{d}x\,\mathrm{d}y=\int_{\mathcal{I}}\int_{\mathcal{I}}b^{N}\left(\ell^{N}\right)^{\frac{\gamma+1}{\gamma}}w^{N}\varphi\,\mathrm{d}x\,\mathrm{d}y.$
We write
$\int_{\mathcal{I}}\int_{\mathcal{I}}b^{N}\left(\ell^{N}\right)^{\frac{\gamma+1}{\gamma}}w^{N}\varphi\,\mathrm{d}x\,\mathrm{d}y$ $=\int_{\mathcal{I}}\int_{\mathcal{I}}b^{N}\ell^{\frac{\gamma+1}{\gamma}}w\varphi\,\mathrm{d}x\,\mathrm{d}y$
@@
$\quad+\int_{\mathcal{I}}\int_{\mathcal{I}}b^{N}\Big(\left(\ell^{N}\right)^{\frac{\gamma+1}{\gamma}}-\ell^{\frac{\gamma+1}{\gamma}}\Big)w^{N}\varphi\,\mathrm{d}x\,\mathrm{d}y$
$\quad+\int_{\mathcal{I}}\int_{\mathcal{I}}b^{N}\ell^{\frac{\gamma+1}{\gamma}}(w^{N}-w)\varphi\,\mathrm{d}x\,\mathrm{d}y.$
Since $\ell^{N}\to\ell$ in $L^{1}(\mathcal{I}\times\mathcal{I})$ by assumption (3.5), and $\left\|\ell^{N}\right\|_{L^{\infty}(\mathcal{I}\times\mathcal{I})}\leq 1$, the map $t\mapsto t^{\frac{\gamma+1}{\gamma}}$ is Lipschitz continuous on $[0,1]$, and thus $\left(\ell^{N}\right)^{\frac{\gamma+1}{\gamma}}\to\ell^{\frac{\gamma+1}{\gamma}}$ in $L^{1}(\mathcal{I}\times\mathcal{I})$. Moreover, $w^{N}\to w$ in $L^{1}(\mathcal{I}\times\mathcal{I})$ and $\left\|w^{N}\right\|_{L^{\infty}(\mathcal{I}\times\mathcal{I})}\leq 1$. Therefore,
$\Big\|\Big(\left(\ell^{N}\right)^{\frac{\gamma+1}{\gamma}}-\ell^{\frac{\gamma+1}{\gamma}}\Big)w^{N}\Big\|_{L^{2}(\mathcal{I}\times\mathcal{I})}\to 0,\qquad\|w^{N}-w\|_{L^{2}(\mathcal{I}\times\mathcal{I})}\to 0,$
as $N\to\infty$. Using the uniform boundedness of $b^{N}$ in $L^{2}(\mathcal{I}\times\mathcal{I})$, we conclude
$\lim_{N\to\infty}\int_{\mathcal{I}}\int_{\mathcal{I}}b^{N}\left(\ell^{N}\right)^{\frac{\gamma+1}{\gamma}}w^{N}\varphi\,\mathrm{d}x\,\mathrm{d}y$ $=$ $\lim_{N\to\infty}b^{N}\ell^{\frac{\gamma+1}{\gamma}}w\varphi\,\mathrm{d}x\,\mathrm{d}y$
$=$ $\int_{\mathcal{I}}\int_{\mathcal{I}}b\ell^{\frac{\gamma+1}{\gamma}}w\varphi\,\mathrm{d}x\,\mathrm{d}y,$
where we used the weak convergence $b^{N}\rightharpoonup b$ in $L^{2}(\mathcal{I}\times\mathcal{I})$. We thus have $a=b\ell^{\frac{\gamma+1}{\gamma}}w$, and, consequently,
$\begin{split}\mathcal{F}_{\mathrm{met}}[b]&=\frac{\nu}{2\gamma}\iint_{\mathcal{I}\times\mathcal{I}}b(x,y)^{\gamma}\,\ell(x,y)^{\gamma+1}\,\,\mathrm{d}w(x,y)\\ &=\frac{\nu}{2\gamma}\|a\|_{L^{\gamma}(\mathcal{I}\times\mathcal{I})}^{\gamma}\\ &\leq\liminf_{N\to\infty}\mathcal{F}_{\mathrm{met}}^{N}[b^{N}].\end{split}$ (8.2)
@@
and combining this with (8.2) we obtain
@@
We claim that $b$ is a global minimizer of the continuum energy functional (4.4)–(4.5) in $L^{\gamma}_{r}(\mathcal{I}\times\mathcal{I})$. For contradiction, let us assume that there exists $\tilde{b}\in L^{\gamma}_{r}(\mathcal{I}\times\mathcal{I})$ such that
@@
Then Theorem 4.1 asserts that the sequence $\tilde{b}^{N}:=\mathcal{Z}^{N}[\tilde{b}$] converges to $\tilde{b}$ in the norm topology of $L^{\gamma}(\mathcal{I}\times\mathcal{I})$, and (4.10) gives
@@
where $\tilde{B}^{N}\in\mathbb{B}^{N}_{r}$ is the matrix of the values of $\tilde{b}^{N}$, i.e., $\tilde{b}^{N}=\mathcal{Q}^{N}[\tilde{B}^{N}]$. Therefore,
[…]